# Unitary coupled-cluster singles and doubles (UCCSD) calculations in conjunction with fragment molecular orbital (FMO) scheme Kenji Sugisaki , Tatsuya Nakano, Yuji Mochizuki

### 28 Feb 2024 v1

@@ 3.2. Relation with Trotter error
It is interesting to note that the monomers “2” and “3” of (FH)-HO are symmetrically equivalent, but VQE-UCCSD yields different correlation energies. The HF molecular orbitals of the monomers “2” and “3” are illustrated in Figure 2. We found that relative phase of the second and the fourth molecular orbitals is different between monomers “2” and “3”, which causes differences in the absolute sign of excitation amplitudes and . As a result, the quantum state corresponding to the UCCSD wave function is not identical for monomers “2” and “3”, and the Trotter error appears in a different way. This fact is confirmed by performing the UCCSD calculations without Trotterization, using the expm_multiply function in the SciPy library [57], which allows us to compute the action of the matrix exponential of on . In this case, the calculated correlation energies of the monomers “2” and “3” are exactly the same: Hartree.
Since the Trotter errors appears in an unexpected way, we further investigated about the relationship between Trotter errors and size-consistency, which is an essential condition in the FMO framework as mentioned earlier. Here we focused on the tetrahydrogen (4H) cluster [58] in a square coordinate with R(H–H) = 1.0583 Å (2.0 Bohr) as the monomer, because the Trotter error becomes more significant when the HF is not a good approximation of the ground-state wave function and the UCCSD wave function has large excitation amplitudes. This system is also suitable because the HF molecular orbitals are completely defined by point-group symmetry. In the dimer (8H) calculations, two 4H clusters were placed to form a cuboid, with the inter-monomer distance being 100 Å. Two types of molecular orbitals are examined in the dimer calculations: Completely delocalized HF canonical orbitals in D point group (denoted as CMO) and orbitals localized on the monomers (denoted as LMO). In the total energy calculations using UCCSD, we used cuQuantum-based quantum circuit simulations with Trotter decomposition and Trotter-free calculations using the expm_multiply function in SciPy. The results are summarized in Table 5.
| System | Trotter decomposition | |
@@
From Table 5, the UCCSD calculations without Trotter decomposition yield almost the same values for both LMO and CMO-based calculations, and the values of the dimer are twice of that of the monomer; i.e., Trotter-free UCCSD satisfies the size-consistency condition. Small differences in the values of dimer with CMO and LMO are due to rounding errors in the AO MO transformation. In contrast, when the Trotter decomposition is used to construct the UCCSD quantum circuit, the of dimer with CMO is significantly larger than the . Since the value for dimer with LMO is approximately double of the of monomer, we concluded that size-consistency condition of the VQE-UCCSD can be maintained when the molecular orbitals localized on each monomer are used in the calculations. This observation is very important for the application of the VQE-UCCSD method to large systems. In the present FMO-UCCSD calculations the molecular orbitals used for the dimers are not completely localized on the monomers, and the size-consistency condition may be broken to some extent. We expect that the use of localized molecular orbitals will improve the correlation energy of the dimers and the sum of the correlation energies, and the FMO-UCCSD calculations with the localized orbitals are in progress. It should be noted that the dependence of the Trotter error on the locality of the molecular orbitals was investigated by Babbush and coworkers [59], and they reported that the Trotter error is larger for localized orbital basis than for canonical orbitals and natural orbitals. The reported study focused only on single molecule (monomer), and our discussions are based on comparing the energies of monomers and a dimer. Our results do not contradict the previous study.
@@ 4. Summary
We have performed VQE-UCCSD calculations [9, 10, 11, 12, 13] using the cuQuantum simulator [18] in conjunction with the FMO calculations for the (FH) and (FH)-HO systems. The STO-3G minimal basis set [46] was used and the frozen-core restriction [47] was imposed. By combining with symmetry conserving Bravyi–Kitaev transformation [51], we can simulate the HFHO dimer with 20 qubits. Both COBYLA and Powell methods were used for VQE optimization, with the latter usually providing better energies but requires more function calls. (FH)-HO system has two symmetrically equivalent FH molecules that should have the same energies, but the Trotterized UCCSD does not satisfy symmetry and yields different energies. The difference in correlation energies between the dimers “21” and “31” is on the order of 1 kcal mol. Size-consistency of the Trotterized UCCSD is also studied numerically using 4H and 8H model clusters. We found that the size-consistency condition can be broken when the molecular orbitals delocalized to the dimer are used for the calculation, and using molecular orbitals localized to the monomers is essential to satisfy the size-consistency. These findings on the relationship between size-consistency and orbital-invariance [31, 39] and the error in the Trotter decomposition are very important not only for FMO-based quantum chemical calculations and other fragmentation-oriented methods [21, 22, 23] but also for VQE-UCCSD in general. Possible future countermeasures include choosing an appropriate optimization algorithm, setting a stricter convergence threshold, using finer Trotter slices, or adopting higher-order Trotter decomposition.
The GPU acceleration was found to be 7.49 and 7.33 with COBYLA and Powell algorithms, respectively, for the dimer “32” of (FH). The usefulness of GPUs has attracted much attention in various fields of quantum computation [60], and quantum chemistry is an example where the acceleration effect is significant [61], including in this case; even with GPU acceleration, it still takes orders of magnitude longer than a regular FMO-CCSD(T) calculation [41]. Recently, an example of large-scale quantum computation with adamantanes has been reported using VQE [62]. Following these trends, we will perform larger FMO-UCCSD computations on upcoming GPU environments.

### 13 Mar 2024 v2

@@ 3.2. Relation with Trotter error
It is interesting to note that the monomers “2” and “3” of (FH)-HO are symmetrically equivalent, but VQE-UCCSD yields different correlation energies. The HF canonical orbitals of the monomers “2” and “3” are illustrated in FIG. 2. We found that the relative phase from the second to the fifth molecular orbitals is different (or inverted) between monomers “2” and “3”, which causes changes in the absolute sign of the some excitation amplitudes and . As a result, the quantum states corresponding to the UCCSD wave function are not identical for monomers “2” and “3”, and the Trotter error appears in a different way. This fact is confirmed by performing the UCCSD calculations without Trotterization, using the expm_multiply function in the SciPy library [68], which allows us to compute the action of the matrix exponential of on . In this case, the calculated correlation energies of the monomers “2” and “3” are exactly the same: Hartree. The fact that Trotterized UCCSD can not maintain orbital-invariance indicates that care must be taken to ensure that the relative phases of the molecular orbitals match at all points when calculating potential energy surfaces.
Since the Trotter errors appears in an unexpected way, we further investigated about the relationship between Trotter errors and the size-consistency, which is an essential condition in the FMO framework as mentioned earlier. Here we focused on the tetrahydrogen (4H) cluster [69] in a square coordinate with R(H–H) = 1.0583 Å (2.0 Bohr) as the monomer, because the Trotter error becomes more significant when the HF is not a good approximation of the ground-state wave function and the UCCSD wave function has large excitation amplitudes. This system is also suitable because the HF CMOs are completely defined by point-group symmetry. In the dimer (8H) calculations, two 4H clusters were placed to form a cuboid, with the inter-monomer distance being 100 Å. Two types of molecular orbitals are examined in the dimer calculations: Completely delocalized canonical orbitals by HF in D point group and LMOs on the monomers. In the total energy calculations using UCCSD:PW, we used cuQuantum-based quantum circuit simulations with Trotter decomposition and without Trotter decomposition using the expm_multiply function in SciPy. The results are summarized in TABLE 5.
| Unit | Trotter decomposition | |
@@
From TABLE 5, the UCCSD calculations without Trotter decomposition yield almost the same values for both LMO- and CMO-based calculations, and the values of the dimer are twice those of the monomer; i.e., Trotter-free UCCSD satisfies the size-consistency condition. Small differences in the values of the dimer with CMO and LMO are due to rounding errors in the AO MO transformation. In contrast, when the Trotter decomposition is used to construct the UCCSD quantum circuit, the of the dimer with CMO is significantly larger than the . Since the value for dimer with LMO is approximately twice the of monomer, we concluded that the size-consistency condition of the VQE-UCCSD can be maintained when the molecular orbitals localized on each monomer are used in the calculations.
In the present study, we used the first-order Trotter decomposition given in Eq. (5), with the number of Trotter slices . To further investigate the relationship between Trotter error and the size-consistency, we run the UCCSD:PW simulations with changed from 1 to 5. We also carried out the UCCSD:PW simulations with the second-order Trotter decomposition given in Eq. (6).
The results are summarized in TABLE S3 in the Supplementary Materials. The simulations ended within one hour when GPU is used. The UCCSD:PW energies of monomer (4H cluster) and dimer (8H cluster) with LMO do not change by using a larger number of Trotter slices or by adopting the second-order Trotter decomposition. For the dimer calculations with CMO, in contrast, the deviation from the CAS-CI energy systematically decreases with increasing . However, even for , the UCCSD:PW energy does not converge to the Trotter-free UCCSD energy calculated by using expm_multiply in SciPy. This result also implies the importance of using LMO in the FMO scheme in conjunction with the VQE-UCCSD.
Since size-consistency is pivotal not only for FMO but also for general quantum chemical calculations, it is important to provide methods to estimate the Trotter error-free energy. Here we examined the extrapolation method to infer the Trotter error-free UCCSD energy using the idea of algorithmic error mitigation [70]. To do this, we plotted the UCCSD:PW energies as a function of the inverse of the number of Trotter slices, , and fitted with a function . It should be noted that in the context of Hamiltonian simulations, the error of the first-order Trotter decomposition scales as [71]. In VQE, however, different Trotterized versions of the UCCSD correspond to different ansatzes, and thus the optimal variational parameters are different. Therefore, it is not necessary to scale the Trotter error of UCCSD as . The calculation results are illustrated in FIG. 3. The E(UCCSD:PW) were successfully fitted by the function , and the difference between the energies estimated from the extrapolation and the Trotter error-free one calculated with expm_multiply is only 35 Hartree. We expect that the extrapolation method used in this work will help to obtain the VQE-UCCSD energy with the size-consistency condition.
It should be noted that the dependence of the Trotter error on the locality of the molecular orbitals was investigated by Babbush and coworkers [72], and they reported that the Trotter error is larger for localized orbital basis than for canonical orbitals and natural orbitals. The reported study focused only on single molecule (monomer), and our discussions are based on comparing the energies of monomers and a dimer. Our results do not contradict this previous study.
@@ 4. Summary
We have performed the VQE-UCCSD calculations [16, 17, 18, 19, 20] using the cuQuantum simulator [25] in conjunction with the FMO calculations for the (FH) and (FH)-HO systems. The STO-3G minimal basis set [56] was used and the frozen-core restriction [57] was imposed. By combining with symmetry conserving Bravyi–Kitaev transformation [62], we can simulate the HFHO dimer with 20 qubits. Both COBYLA and Powell methods were used for VQE optimization, with the latter usually providing better energies but requires more function calls. When the HF CMOs were used for the UCCSD wave function expansion, the calculated correlation energies were somewhat small in magnitude, possibly due to the breakdown of the size-consistency condition. By using the LMOs, the UCCSD correlation energies improved dramatically, and the differences in the correlation energies between the CAS-CI and the UCCSD in conjunction with the Powell optimizer were calculated to be 0.71 and 1.29 kcal mol for (FH) and (FH)-HO, respectively. The (FH)-HO system has two symmetrically equivalent FH molecules that should have the same energies, but the Trotterized UCCSD does not satisfy symmetry and yields different energies. The difference in correlation energies between the dimers “21” and “31” is on the order of 1 kcal mol in the canonical orbital basis but it reduced to 0.039 kcal mol in the localized orbital basis. Size-consistency of the Trotterized UCCSD is also studied numerically using 4H and 8H model clusters. We found that the size-consistency condition can be broken when the molecular orbitals delocalized to the dimer are used for the calculation, and using molecular orbitals localized to the monomers is essential to satisfy the size-consistency. These findings on the relationship between size-consistency and orbital-invariance [39, 49] and the error in the Trotter decomposition are very important not only for FMO-based quantum chemical calculations and other fragmentation-oriented methods [28, 29, 30] but also for VQE-UCCSD in general. From the numerical simulations, we also demonstrated that the Trotter error-free UCCSD energy can be estimated by means of extrapolation by computing the UCCSD energies with different numbers of Trotter slices.
The GPU acceleration was found to be 7.30 and 7.64 with COBYLA and Powell algorithms, respectively, for the dimer “21” of (FH) (18 qubit system). For the dimer “21” of (FH)-HO, the estimated GPU acceleration ratio of the VQE quantum circuit simulation to be about 30.85, and the VQE simulations of the dimer “21” with COBYLA and Powell will take about 21 and 54 days, respectively. The usefulness of GPUs has attracted much attention in various fields of quantum computation [73], and quantum chemistry is an example where the acceleration effect is significant [74], including in this case; even with GPU acceleration, it still takes orders of magnitude longer than a regular FMO-CCSD(T) calculation [51]. Recently, an example of large-scale quantum computation with adamantanes has been reported using VQE [75]. Following these trends, we will perform larger FMO-UCCSD computations on upcoming GPU environments.
@@ 5. Acknowledgement
Supplementary Materials
| Unit | CMO | LMO | ||
| COBYLA | Powell | COBYLA | Powell | |
| Monomer | ||||
| “1” | 115 | 225 | 133 | 256 |
| “2” | 130 | 257 | 113 | 257 |
| “3” | 119 | 193 | 125 | 241 |
| Dimer | ||||
| “21” | 1762 | 4783 | 1924 | 5264 |
| “31” | 1963 | 4765 | 2085 | 5188 |
| “32” | 1910 | 5408 | 1546 | 5188 |
| Unit | CMO | LMO | ||
| COBYLA | Powell | COBYLA | Powell | |
| Monomer | ||||
| “1” | 262 | 401 | 474 | 791 |
| “2” | 114 | 238 | 142 | 246 |
| “3” | 121 | 215 | 129 | 207 |
| Dimer | ||||
| “21” | 3278 | 12448 | 4913 | 12452 |
| “31” | 3511 | 11966 | 4568 | 12764 |
| “32” | 1348 | 2544 | 1342 | 3882 |
| Unit | First-order Trotter | Second-order Trotter | expm_multiply111111Trotter-free implementation using expm_multiply in SciPy library. | ||||
| 121212 represents the number of Trotter slices. | |||||||
| Monomer | 0.8102 | 0.8101 | 0.8099 | 0.8100 | 0.8100 | 0.8101 | 0.8118 |
| Dimer (LMO) | 1.6207 | 1.6200 | 1.6203 | 1.6204 | 1.6205 | 1.6206 | 1.6236 |
| Dimer (CMO) | 5.0319 | 1.9139 | 1.7141 | 1.6681 | 1.6496 | 2.0172 | 1.6234 |